solve this proportion for the variable: x+2/x+6 = x-4/x-2
A) 7
B) -7
C) -10
D) 10

Answer by stanbon(48558)

solve this proportion for the variable: x+2/x+6 = x-4/x-2
(x+2)(x-2) = (x+6)(x-4)
x^2-4 = x^2+2x-24

-4 = 2x-24
x -12 = -2
x = 10
===========
Cheers,
Stan H.






A) 7
B) -7
C) -10
D) 10